Compute the price index for each year. Use the first year as the base year. What was the inflation rate between the two years. Item Quantity Unit Price-Last Year Unit Price-This Year Coffee 20 pounds $3.00 $4.00 Tuition 1 year 4,000.00 7,000.00 Pizza 100 pizzas 8.00 10.00 VCR rental 75 days 15.00 10.00 Vacation 2 weeks 300.00 500.00


1
Expert's answer
2021-04-28T09:46:01-0400

Price index = Cost of basket in current year/Cost of basket in base year × 100


Price index (Last Year)


= [20 × 3 + 1 × 4000 + 100 § 8 + 75 × 15 + 2 × 300]/[20 × 3 + 1 x 4000 + 100 § 8 + 75 § 15 + 2 × 300] × 100


= [60 + 4000 + 800 + 1125 + 600]/6585 × 100


= 6585/6585 × 100 = 100


Price index (This year)


= [20 × 4 + 1 × 7000 + 100 × 10 + 75 × 10 + 2 × 500]/[20 × 3 + 1 × 4000 + 100 × 8 + 75 × 15 + 2 × 300] ×100


= 9830/6585 × 100 = 149.28


Inflation = (Price index in this year "-" Price index in base year)/Price index in base year ×100


= (149.28 "-" 100)/100 ×100 = 49.28%